40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 7 11:25:31 2020, gautam, Update, SUS, MC1 suspension glitchy again 
    Reply  Wed Dec 9 16:22:57 2020, gautam, Update, SUS, Yet another round of Sat. Box. switcharoo WFS2.pngWFS_lineNoise.pngWFSchar.pdf
       Reply  Sun Jan 3 16:26:06 2021, Koji, Update, SUS, IMC WFS check (Yet another round of Sat. Box. switcharoo) Screen_Shot_2021-01-03_at_17.14.57.png
    Reply  Thu May 13 11:55:04 2021, Anchal, Paco, Update, SUS, MC1 suspension misbehaving Screenshot_from_2021-05-13_09-50-24.pngMC1_Glitches_Invest2.pdf
       Reply  Thu May 13 19:38:54 2021, Anchal, Update, SUS, MC1 Satellite Amplifier Debugged MC1_UL_Channel_Fixed.png
          Reply  Mon May 24 19:14:15 2021, Anchal, Paco, Summary, SUS, MC1 Free Swing Test set to trigger 
             Reply  Tue May 25 10:22:16 2021, Anchal, Paco, Summary, SUS, MC1 new input matrix calculated and uploaded SUS_Input_Matrix_Diagonalization.pdf
          Reply  Thu Jun 17 11:45:42 2021, Anchal, Paco, Update, SUS, MC1 Gave trouble again SummaryScreenShot.pngMC1_LL_SENSOR_DEAD.png
             Reply  Thu Jun 17 16:37:23 2021, Anchal, Paco, Update, SUS, c1susaux computer rebooted 
                Reply  Tue Jun 22 11:56:16 2021, Anchal, Paco, Update, SUS, ADC/Slow channels issues CDS_FE_Status.png
                   Reply  Tue Jun 22 16:52:28 2021, Paco, Update, SUS, ADC/Slow channels issues shake_and_damp.png
                      Reply  Wed Jun 23 09:05:02 2021, Anchal, Update, SUS, MC lock acquired back again 
                         Reply  Thu Jun 24 16:40:37 2021, Koji, Update, SUS, MC lock acquired back again P_20210624_163641_1.jpg
Message ID: 15720     Entry time: Wed Dec 9 16:22:57 2020     In reply to: 15712     Reply to this: 15747
Author: gautam 
Type: Update 
Category: SUS 
Subject: Yet another round of Sat. Box. switcharoo 

As discussed at the meeting, I decided to effect a satellite box swap for the misbehaving MC1 unit. I looked back at the summary pages Vmon for the SRM channels, and found that in the last month or so, there wasn't any significant evidence of glitchiness. So I decided to effect that swap at ~4pm today. The sequence of steps was:

  • SRM and MC1 watchdogs were disabled.
  • Unplugged the two satellite boxes from the vacuum flanges.
  • For the record: S/N 102 was installed at MC1, and S/N 104 was installed at SRM. Both were de-lidded, supposedly to mitigate the horrible thermal environment a bit. S/N 104 was the one Koji repaired in Aug 2019 (the serial number isn't visible or noted there, but only one box has jumper wires and Koji's photos show the same jumper wires). In June 2020, I found that the repaired box was glitching again, which is when I swapped it for S/N 102.
  • After swapping the two units, I re-enabled the local damping on both optics, and was able to re-lock the IMC no issues.

One thing I was reminded of is that the motion of the MC1 optic by controlling the bias sliders is highly cross-coupled in pitch and yaw, it is almost diagonal. If this is true for the fast actuation path too, that's not great. I didn't check it just now.

While I was working on this, I took the opportunity to also check the functionality of the RF path of the IMC WFS. Both WFS heads seem to now respond to angular motion of the IMC mirror - I once again dithered MC2 and looked at the demodulated signals, and see variation at the dither frequency, see Attachment #1. However, the signals seem highly polluted with strong 60 Hz and harmonics, see the zoomed-in time domain trace in Attachment #2. This should be fixed. Also, the WFS loop needs some re-tuning. In the current config, it actually makes the MC2T RIN worse, see Attachment #3 (reference traces are with WFS loop enabled, live traces are with the loop disabled - sorry for the confusing notation, I overwrote the patched version of DTT that I got from Erik that allows the user legend feature, working on getting that back).

Quote:

The MC1 suspension has begun to show evidence of glitches again, from Friday/Saturday. You can look at the suspension Vmon tab a few days ago and see that the excess fuzz in the Vmon was not there before. The extra motion is also clearly evident on the MCREFL spot. I noticed this on Saturday evening as I was trying to recover the IMC locking, but I thought it might be Millikan so I didn't look into it further. Usually this is symptomatic of some Satellite box issues. I am not going to attempt to debug this anymore.

Attachment 1: WFS2.png  404 kB  | Hide | Hide all
WFS2.png
Attachment 2: WFS_lineNoise.png  201 kB  | Hide | Hide all
WFS_lineNoise.png
Attachment 3: WFSchar.pdf  109 kB  | Hide | Hide all
WFSchar.pdf
ELOG V3.1.3-